# Model Provenance via Model DNA

### v2

@@ 2 Related work
Representation Learning. Representation learning refers to the process of learning a parametric mapping from the raw input data domain to a feature vector, in the hope of capturing and extracting more abstract and useful concepts that can improve performance on a range of downstream tasks. The aim of learning representations of the data is to make it easier to extract useful information when building classifiers or other predictors. Representation learning has played a tremendous role in different SOTA frameworks [2]. Recently, Contrastive Representation Learning (CRL) is widely used in NLP and CV [17]. It can be considered as self-supervised learning by comparing among different samples [4, 33]. In this paper, we explore a way of machine learning model representation. Our approach involves a data-driven and model-driven representation learning framework that constructs a model representation in a latent space. Additionally, there are some works to investigate the relationship of Neural Network representations [6, 14]. This work is novel since it is the first time, as far as we know, that such an approach is proposed for representing the ML model.
@@ 3.1 Definition
Consider a dataset $D_{s}=\{(x_{i},y_{i})\}_{i=1}^{{|D_{s}|}}$ where each $x_{i}\in\mathbb{R}^{d}$ is a data instance and $y_{i}\in Y=\{1,2,\ldots,c\}$ is its associated class label. Using an algorithm architecture $\mathcal{A}_{s}$, a machine learning model $M_{s}$ is learned from $D_{s}$. Let $D_{t}={(x_{j},y_{j})}_{j=1}^{{|D_{t}|}}$ be a dataset used to learn another machine learning model $M_{t}$. If $M_{t}$ is learned using the pre-training model $M_{s}$ with the same algorithm architecture $\mathcal{A}_{s}$, we refer to $M_{s}$ as the source model and $M_{t}$ as the homologous model of $M_{s}$. Similarly, let $\bar{M}_{t}$ be a machine learning model learned from the dataset $D_{t}={(x_{j},y_{j})}_{j=1}^{{|D_{t}|}}$ but based on random initialization or pre-training without using $M_{s}$. We refer $\bar{M}_{t}$ as non-homologous to $M_{s}$.
The problem of Model Provenance (MP) is to find a function $f$ such that for a given source model $M_{s}$ and any target model $M_{1}$,
$f(M_{s},M_{1})=\begin{cases}1,&\mbox{if $M_{1}$ is the homologous model of $M_{s}$.}\\ 0,&\mbox{otherwise}.\end{cases}$ (1)
When $f(M_{s},M_{1})=1$, we say $M_{s}$ is the provenance of $M_{1}$.
@@ 3.2 Discussion
Furthermore, we conducted experiments using the standard continual learning setup on the image classification task[32, 24] to show the relationship between homologous models. Specifically, we examined two model architectures, ResNet18 [9] and AlexNet [15], on the CIFAR10 and CIFAR100 datasets. A source model $M_{s}$ is first trained on CIFAR10. Then we randomly select ten classes from CIFAR100 to train target models $M_{t}$. Note that $M_{t}$ is learned by initializing its parameters using those of the source model $M_{s}$. In Figure 1, we present the results of the experiments conducted on ResNet18, and the results for AlexNet can be found in the Appendix. The background color indicates the different training phases. The red color represents the training of source model $M_{s}$ on the CIFAR10 dataset, while the blue color represents the training of target model $M_{t}$ on the CIFAR100 dataset. The red line corresponds to the accuracy of testing the model on the training data of CIFAR10, while the blue line represents the accuracy of testing the model on the training data of CIFAR100. In the blue background part, we train the model based on the parameters of the model obtained in the red background part.
@@
Figure 1 (a) demonstrates that the predictive accuracy of $M_{t}$ is significantly reduced when evaluated on the CIFAR10 dataset. The red section represents a model (called source model) trained on the source domain (CIFAR10), while the blue section represents this model (called target model) continuously trained on the target domain (CIFAR100). The red line indicates model evaluation on CIFAR10, and the blue line represents model evaluation on CIFAR100. This observation suggests that the training data of the source model may have been severely forgotten in the homologous model (i.e., large $||\Delta w||$). Figures 1 (b)-(d) provide further insight into the relationship between the source model $M_{s}$ and the homologous model $M_{t}$. They show the predictive accuracy of $M_{t}$ on the CIFAR10 dataset when different numbers of the last layers of $M_{t}$ are replaced with corresponding layers in $M_{s}$. In Figure. (b), within the blue section, we replace the model’s last layer with the model from the red section. Consequently, the model exhibits reduced performance when evaluated on CIFAR10 (red line). Similarly, in Figures. (c) and (d), we extended this by replacing the model’s last two/three layers with the ones from the red section, respectively. We can observe that as we replace more layers in $M_{t}$ with those from $M_{s}$, the predictive accuracy of $M_{t}$ on the CIFAR10 dataset increases. In Figure. (e), we consider a scenario where in the blue section, the (target) model is not initialized by the (source) model from the red section. Instead, the (target) model is trained with random initialization. In this setting, we replace the model’s last three layers with the corresponding layers from the (source) model in the red section (as depicted in Figure. (a)). Figure 1 (e) shows that even with the last layers replaced, the predictive accuracy of $\bar{M}_{t}$ is still poor. The green line in Figure 1 (e) represents the training accuracy of $\bar{M}_{t}$. Figures. (b) and (e) illustrate the presence of a relationship between the source model and its homologous model. However, non-homologous models, do not exhibit any discernible relationship even by replacing the layers between them.
@@ 4.2 MGMP
We introduce a framework termed Model DNA Generation and Model Provenance identification (MGMP), which operates on the principles of deep representation learning. The MGMP framework comprises three primary components: DNA generation, DNA similarity loss, and provenance classifier, as depicted in Figure 2. The process commences with the source model’s training data, which is fed into the DNA generator alongside the source and target models. This concatenation produces outputs from the DNA generator, source model, and target model, which are then amalgamated to generate the model DNA, exemplified by the integration of DNA generator outcomes and model predictions. Subsequently, we formulate the model DNA for both the source and target models. To preserve information relevant to homologous and non-homologous models, a DNA similarity loss is incorporated. Finally, we exploit the model DNA to train a binary prediction network, allowing us to infer provenance outcomes. Each component of the MGMP framework is described in subsequent sections.
@@ 4.2.2 DNA similarity loss
The DNA similarity loss is designed with a specific purpose: to measure the similarity between the DNA of two models and to satisfy the properties of DNA latent space. We achieve this by employing a metric that ensures that the similarity of $(o^{s}_{i},o^{t}_{i})$ in the latent space is not disclosed to $(o^{s}_{i},\bar{o}^{t}_{i})$, which is inspired by Contrastive Learning [4, 35]. In more detail, our approach doesn’t just preserve the relationships like $(o^{s}_{i},o^{t}_{i})$ and $(o^{s}_{i},\bar{o}^{t}_{i})$ based on a single input $x_{i}$. Instead, it takes into account the amalgamation of information across relationships present in $\mathbb{O}_{s}$, $\mathbb{O}_{t}$, and $\mathbb{\bar{O}}_{t}$. This implies that we incorporate more relationships into the loss function, depicted in Figure 3. The solid line represents relationships used in prior works, while the dashed line signifies the relationship between distinct DNA fragments within a model.
@@
Remark: In practice, after assembling the DNA fragments, we consider a cosine distance in our final decision (Eqs. (5) and (6)). Cosine similarity is the cosine of the angle between the vectors; that is, it is the dot product of the vectors divided by the product of their lengths. It follows that the cosine similarity does not depend on the magnitudes of the vectors, but only on their angle. Therefore, through DNA assemble with addition, the angles between the representations of the source and target models change even though the vectors only linearly move. That is why the generator can improve the results. Meanwhile, we emphasize that the part of DNA assemble is flexible, which means it can be replaced by different combining methods, e.g., concatenation or adding some networks (e.g., fully connected layers).
@@ 4.2.3 Provenance classifier
[…]
@@ 4.2.5 Prediction
In the prediction phase, our framework is equipped to provide predictions at various levels of granularity concerning the data. Specifically, we can choose any model (e.g., $M_{1}$) from the target model set to ascertain its homology with the source model $M_{s}$, based on either the DNA fragment $o$ or the DNA set $\mathbb{O}$. For predictions at the level of DNA fragments, the process is carried out as follows:
$f(x_{i},M_{s},M_{1})=\begin{cases}1,&\text{if}\ f_{p}(o^{c}_{i})\rightarrow 1\\ 0,&\text{otherwise}\end{cases}$ (9)
where $o^{c}_{i}$ is the concatenation of DNA fragment $o^{s}_{i}$ and $o^{1}_{i}$ corresponding to $x_{i}$ for $x_{i}\in D_{s}$. For the granularity of the model DNA set, we can compute the average results by considering the whole DNA fragments: $f(D_{s},M_{s},M_{1})=1$ if $\frac{1}{N}\sum_{i=1}^{N}f_{p}(o^{c}_{i})\rightarrow 1$, and 00 otherwise [^3] In practice, we can also employ a threshold $\delta$ for determining the final prediction, i.e., $\frac{1}{N}\sum_{i=1}^{N}f_{p}(o^{c}_{i})\geq\delta$. and $\delta$ can be seen as a degree of similarity of DNA. The choice of threshold depends on the specific context and requirements of the problem at hand..

### v3

@@ 2 Related work
Representation Learning. Representation learning refers to the process of learning a parametric mapping from the raw input data domain to a feature vector, in the hope of capturing and extracting more abstract and useful concepts that can improve performance on a range of downstream tasks. The aim of learning representations of the data is to make it easier to extract useful information when building classifiers or other predictors. Representation learning has played a tremendous role in different frameworks [2]. Recently, Contrastive Representation Learning (CRL) is widely used in NLP and CV [17]. It can be considered as self-supervised learning by comparing among different samples [4, 33]. In this paper, we explore a way of machine learning model representation. Our approach involves a data-driven and model-driven representation learning framework that constructs a model representation in a latent space. Additionally, there are some works to investigate the relationship of Neural Network representations [6, 14]. This work is novel since it is the first time, as far as we know, that such an approach is proposed for representing the ML model.
@@ 3.1 Definition
Consider a dataset $D_{s}=\{(x_{i},y_{i})\}_{i=1}^{{|D_{s}|}}$ where each $x_{i}\in\mathbb{R}^{d}$ is a data instance and $y_{i}\in Y=\{1,2,\ldots,c\}$ is its associated class label, a machine learning model $M_{s}$ is learned from $D_{s}$ by using an algorithm architecture $\mathcal{A}_{s}$. Let $D_{t}=\{(x_{j},y_{j})\}_{j=1}^{{|D_{t}|}}$ be a dataset used to learn a machine learning model $M_{t}$. If $M_{t}$ is learned using the pre-training model $M_{s}$ with the same algorithm architecture $\mathcal{A}_{s}$, we refer to $M_{s}$ as the source model and $M_{t}$ as the homologous model of $M_{s}$. Similarly, let $\bar{M}_{t}$ be a machine learning model learned from the dataset $D_{t}={(x_{j},y_{j})}_{j=1}^{{|D_{t}|}}$ but based on random initialization or pre-training without using $M_{s}$. We refer $\bar{M}_{t}$ as non-homologous to $M_{s}$.
The problem of Model Provenance (MP) is to find a function $f$ such that for a given source model $M_{s}$ and any target model $M_{t}$,
$f(M_{s},M_{t})=\begin{cases}1,&\mbox{if $M_{t}$ is the homologous model of $M_{s}$.}\\ 0,&\mbox{otherwise}.\end{cases}$ (1)
When $f(M_{s},M_{t})=1$, we say $M_{s}$ is the provenance of $M_{t}$.
@@ 3.2 Discussion
Furthermore, we conducted experiments using the standard continual learning setup on the image classification task [32, 24] to show the relationship between homologous models. Specifically, we examined two model architectures, ResNet18 [9] and AlexNet [15], on the CIFAR10 and CIFAR100 datasets. A source model $M_{s}$ is first trained on CIFAR10. Then we randomly select ten classes from CIFAR100 to train target models $M_{t}$. Note that $M_{t}$ is learned by initializing its parameters using those of the source model $M_{s}$. In Figure 1, we present the results of the experiments conducted on ResNet18, and the results for AlexNet can be found in [mu2023model]. The background color indicates the different training phases. The red color represents the training of source model $M_{s}$ on the CIFAR10 dataset, while the blue color represents the training of target model $M_{t}$ on the CIFAR100 dataset. The red line corresponds to the accuracy of testing the model on the training data of CIFAR10, while the blue line represents the accuracy of testing the model on the training data of CIFAR100. In the blue background part, we train the model based on the parameters of the model obtained in the red background part.
@@
Figure 1 (a) demonstrates that the predictive accuracy of $M_{t}$ is significantly reduced when evaluated on the CIFAR10 dataset. The red section represents a model (called source model) trained on the source domain (CIFAR10), while the blue section represents this model (called target model) continuously trained on the target domain (CIFAR100). The red line indicates model evaluation on CIFAR10, and the blue line represents model evaluation on CIFAR100. This observation suggests that the training data of the source model may have been severely forgotten in the homologous model (i.e., large $||\Delta w||$). Figures 1 (b)-(d) provide further insight into the relationship between the source model $M_{s}$ and the homologous model $M_{t}$. They show the predictive accuracy of $M_{t}$ on the CIFAR10 dataset when different numbers of the last layers of $M_{t}$ are replaced with corresponding layers in $M_{s}$. In Figure 1 (b), within the blue section, we replace the model’s last layer with the model from the red section. Consequently, the model exhibits reduced performance when evaluated on CIFAR10 (red line). Similarly, in Figures 1 (c) and (d), we extended this by replacing the model’s last two/three layers with the ones from the red section, respectively. We can observe that as we replace more layers in $M_{t}$ with those from $M_{s}$, the predictive accuracy of $M_{t}$ on the CIFAR10 dataset increases. In Figure 1 (e), we consider a scenario where in the blue section, the (target) model is not initialized by the (source) model from the red section. Instead, the (target) model is trained with random initialization. In this setting, we replace the model’s last three layers with the corresponding layers from the (source) model in the red section (as depicted in Figure 1 (a)). Figure 1 (e) shows that even with the last layers replaced, the predictive accuracy of $\bar{M}_{t}$ is still poor. The green line in Figure 1 (e) represents the training accuracy of $\bar{M}_{t}$. Figures 1 (b) and (e) illustrate the presence of a relationship between the source model and its homologous model. However, non-homologous models, do not exhibit any discernible relationship even by replacing the layers between them.
@@ 4.2 MGMP
We introduce a framework termed Model DNA Generation and Model Provenance identification (MGMP), which operates on the principles of deep representation learning. The MGMP framework comprises three primary components: DNA generation, DNA similarity loss, and provenance classifier, as depicted in Figure 2. The process commences with the source model’s training data, which is fed into the DNA generator alongside the source and target models. This concatenation produces outputs from the DNA generator, source model, and target model, which are then amalgamated to generate the model DNA, exemplified by the integration of DNA generator outcomes and model predictions. Subsequently, we formulate the model DNA for both the source and target models. To preserve information relevant to homologous and non-homologous models, a DNA similarity loss is incorporated. Finally, we exploit the model DNA to train a binary prediction network, allowing us to infer provenance outcomes. Each component of the MGMP is described in subsequent sections.
@@ 4.2.2 DNA similarity loss
The DNA similarity loss is designed with a specific purpose: to measure the similarity between the DNA of two models and to satisfy the properties of DNA latent space. We achieve this by employing a metric that ensures that the similarity of $(o^{s}_{i},o^{t}_{i})$ in the latent space is not disclosed to $(o^{s}_{i},\bar{o}^{t}_{i})$, which is inspired by Contrastive Learning [4, 35]. In more detail, our approach does not just preserve the relationships like $(o^{s}_{i},o^{t}_{i})$ and $(o^{s}_{i},\bar{o}^{t}_{i})$ based on a single input $x_{i}$. Instead, it takes into account the amalgamation of information across relationships present in $\mathbb{O}_{s}$, $\mathbb{O}_{t}$, and $\mathbb{\bar{O}}_{t}$. This implies that we incorporate more relationships into the loss function, depicted in Figure 3. The solid line represents relationships used in prior works, while the dashed line signifies the relationship between distinct DNA fragments within a model.
@@
Remark: In practice, after assembling the DNA fragments, we consider a cosine distance in our final decision (Eqn. (5) and (6)). Cosine similarity is the cosine of the angle between the vectors; that is, it is the dot product of the vectors divided by the product of their lengths. It follows that the cosine similarity does not depend on the magnitudes of the vectors, but only on their angle. Therefore, through DNA assemble with addition, the angles between the representations of the source and target models change even though the vectors only linearly move. That is why the generator can improve the results. Meanwhile, we emphasize that the part of DNA assemble is flexible, which means it can be replaced by different combining methods, e.g., concatenation or adding some networks (e.g., fully connected layers).
@@ 4.2.3 Provenance classifier
where $h_{i}$ is the truth label on each input $o_{i}^{c}$ and $h_{i}^{\prime}\leftarrow f_{p}(o_{i}^{c})$. Note that $h=1$ if $M_{s}$ and $M_{t}$ are homologous and 00 otherwise.
@@ 4.2.5 Prediction
In the prediction phase, our framework is equipped to provide predictions at various levels of granularity concerning the data. Specifically, we can choose any model (e.g., $M_{t}$) to ascertain its homology with the source model $M_{s}$, based on either the DNA fragment $o$ or the DNA set $\mathbb{O}$. For predictions at the level of DNA fragments, the process is carried out as follows:
$f(x_{i},M_{s},M_{t})=\begin{cases}1,&\text{if}\ f_{p}(o^{c}_{i})\rightarrow 1\\ 0,&\text{otherwise}\end{cases}$ (9)
where $o^{c}_{i}$ is the concatenation of DNA fragment $o^{s}_{i}$ and $o^{t}_{i}$ corresponding to $x_{i}$ for $x_{i}\in D_{s}$. For the granularity of the model DNA set, we can compute the average results by considering the whole DNA fragments: $f(D_{s},M_{s},M_{t})=1$ if $\frac{1}{N}\sum_{i=1}^{N}f_{p}(o^{c}_{i})\rightarrow 1$, and 00 otherwise. In practice, we can also employ a threshold $\delta$ for determining the final prediction, i.e., $\frac{1}{N}\sum_{i=1}^{N}f_{p}(o^{c}_{i})\geq\delta$. and $\delta$ can be seen as a degree of similarity of DNA. The choice of threshold depends on the specific context and requirements of the problem at hand.